5(t-3)-7t = -13
I've tried: 5t -15 - 7t = -13; 5t - 7t = -13 + 15; 2t = 2; t = 1 but that doesn't seem to work out. Know I'm doing something wrong, but not sure what.

Answer by Alan3354(69443)

5(t-3)-7t = -13
5t-15 - 7t = -13
-2t-15 = -13
-2t = 2
t = -1